OpenStudy (chrisplusian):

how do you integrate 1-(tan(x))^2 ????? HElP

OpenStudy (chrisplusian):

\[\int\limits 1-(\tan(x))^2\]

OpenStudy (zepp):

\[\int 1-tan^2x\]

OpenStudy (chrisplusian):

yes

OpenStudy (chrisplusian):

sorry with respect to x

OpenStudy (zepp):

You can separate it into two integrals: \[\large \int 1dx -\int tan^2(x)dx\]

OpenStudy (chrisplusian):

Hahahahhaha I have been at it toooooo long I should have seen that it is easy!!!!

OpenStudy (chrisplusian):

thanks

OpenStudy (zepp):

Welcome :)

OpenStudy (chrisplusian):

ok @zepp I am now scratching my head because I know how to integrate tanx but not \[\tan ^{2}x\]

OpenStudy (zepp):

Okay, first recall the trig identity: \(\large sin^2(x)+cos^2(x)=1\).

OpenStudy (zepp):

Divide everything by \(cos^2x\): \[\large \frac{sin^2x}{cos^2x}+\frac{cos^2x}{cos^2x}=\frac{1}{cos^2x}\\\large tan^2x+1=sec^2x\\\large tan^x=sec^2x\]

OpenStudy (zepp):

Now we have \[\large \int1-tan^2xdx=\int1dx-\int tan^2xdx\\\large=\int1dx-\int sec^2x-1dx=\int 1dx-\int sec^2xdx-\int 1dx\]For the last line of the post above, it should be \(\large tan^2x=sec^2x-1\), sorry about that mistake

OpenStudy (chrisplusian):

ok thanks that makes sense

OpenStudy (zepp):

np

OpenStudy (zepp):

Need a parantheses here\[\large \int 1dx-(\int sec^2(x)dx-\int1dx)\]So it would be\[\large \int 1dx-\int sec^2(x)dx+\int 1 dx=-tan(x)+2x+C\]